OpenStudy (anonymous):

Please help me solve this problem A line on a graph passes through the points (-1,3) and (5,15). Write the equation for the line.

OpenStudy (anonymous):

y-3 = 15-3 /5+1 (x+1) => y-3 = 2(x+1) => y = 2x +5

OpenStudy (anonymous):

Thanks